OpenStudy (anonymous):

PLEASE HELP ASAP! A lottery winner must decide between two methods of payment. Choice one is to receive a lump sum of $100,000,000. Choice two is to begin with one cent on day 1, five cents on day 2, 25 cents on day three, and so on until the end of 16 days. Part 1: Which choice should the lottery winner select? (4 points) Part 2: Using complete sentences, explain the procedure taken to answer this question. (4 points)

OpenStudy (hoblos):

the second choice is the sum of a geometric sequence of ratio 5 the total payment by the second choice is \[\sum_{n=0}^{16-1}5^{n} = \frac{ 1-5^{16} }{ 1-5 }\]

OpenStudy (anonymous):

I got this as my answer S = 381469726.56

OpenStudy (anonymous):

is that correct

OpenStudy (hoblos):

it's correct so which is bigger $100,000,000 or 381469726.56 cents ?

OpenStudy (anonymous):

second option! but what method did you use?

OpenStudy (anonymous):

Am i correct?

OpenStudy (hoblos):

the sum of a geometric sequence is \[\sum_{i=0}^{n-1}r ^{i} = \frac{ 1-r ^{n} }{ 1-r }\]